Circular No. 6994

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998dn IN NGC 337A L. Cao, on behalf of the Beijing Astronomical Observatory (BAO)

Supernova Survey, reports the discovery of a supernova (mag 15.8) found on an unfiltered CCD frame taken on Aug. 19.76 UT. SN 1998dn is located at α = 1^{h}01^{m}27^{s}.08, δ = −7^{o}36′36′′.7 (equinox 2000.0), which is 102′′ west and 79′′ south of the center of NGC 337A. A similar observation on Aug. 22.79 confirmed the supernova’s existence at almost the same magnitude. The Digital Sky Survey image and a CCD image taken on 1998 Jan. 15 of the same field show no sign of the new star.

L. Cao and J.-Y. Wei, BAO; and W.-D. Li, University of California at Berkeley, report that a CCD spectrogram (range 380–635 nm) of SN 1988dn was obtained by H.-M. Shi (Shanghai Astronomical Observatory) with the BAO 2.16-m telescope on Aug. 23.72 UT. The spectrum shows a red continuum, along with Hβ and He I 587.6-nm/Na D line features, indicating this to be a type-II supernova well after maximum.

SUPERNOVA 1998dl IN NGC 1084 A. V. Filippenko, University of California at Berkeley, reports that

inspection of a CCD spectrum of SN 1998dl (cf. IAUC 6992) obtained on Aug. 23 by C. De Breuck (Lawrence Livermore National Laboratory) with the 3-m Shane reflector at Lick Observatory reveals that the object is a normal type-II supernova, with well-developed P-Cyg hydrogen Balmer profiles. Na D and the Ca II near-infrared triplet are strong, and there are several prominent Fe II lines.

COMET 52P/HARRINGTON-ABELL This comet continues to be unusually bright (cf. IAUC 6975), as in-

dicated by the following total visual magnitude estimates: July 26.02 UT, 11.2 (M. Reszelski, Szamotuly, Poland, 0.25-m reflector); 30.05, 11.6 (K. Sárneczky, Ágasvár, Hungary, 0.44-m reflector); Aug. 5.13, 12.3 (J. M. Trigo, Benicassim, Castellon, Spain, 0.17-m reflector); 15.97, 12.8 (M. Lehký, Hradec Králové, Czech Republic, 0.42-m reflector); 20.00, 12.3 (M. Plsek, Lelekovice, Czech Republic, 0.20-m reflector); 23.13, 11.8 (J. Carvajal, Avila, Spain, 0.32-m reflector).

1998 August 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green